# Random matrix model of Kolmogorov-Zakharov turbulence

### v1

@@ I Introduction
The Kolmogorov concept of turbulence kolm41 ; obukhov assumes emergence of energy flow through an inertial interval from large spacial scales, with an external pumping, to small scales where energy is absorbed by dissipation. The scaling arguments lead to appearance of a power law energy distribution over wave modes for hydrodynamics turbulence kolm41 ; obukhov . This concept was generalized and extended to weak wave turbulence, based on diagrammatic techniques and the kinetic equation, indeed showing the emergence of power law distributions for various types of weakly interacting nonlinear waves filonenko ; zakharovbook ; nazarenkobook ; luka ; galtier . This theory became known as Kolmogorov-Zakharov (KZ) turbulence (or spectra) nazarenkobook ; luka ; galtier . In spite of various successful confirmations of this theory in experiments and numerical simulations(see e.g. nazarenkobook ; luka ; galtier ) it is still based on a fundamental hypothesis directly stated in the seminal work of Zhakharov and Finonenko filonenko : “In the theory of weak turbulence nonlinearity of waves is assumed to be small; this enables us, using the hypothesis of the random nature of the phase of individual waves, to obtain the kinetic equation for the mean square of the wave amplitudes”. Nevertheless, the dynamical equations for waves do not involve Random Phase Approximation (RPA) and hence, the validity for the whole concept of energy flow from large to small scales remains open.
Indeed, a flow through an inertial interval is Hamiltonian and it is well known that in nonlinear systems with weak nonlinearity the Kolmogorov-Arnold-Moser (KAM) theory guarantees that the main part of the system phase space remains integrable and non-chaotic in the limit of very weak nonlinearity (see e.g. arnold ; sinai ). More physical analysis of nonlinear dynamical Hamiltonian systems also shows the existence of a chaos border below which the phase space contains mainly an integrable dynamics opposite to the turbulent one chirikov ; lichtenberg . Also a simple observation of a sea surface clearly shows that a weak wind is not able to generate turbulence.
@@ II Model description
Here $H_{n,n^{\prime}}$ are elements of an RMT matrix $\hat{H}$ of size $N$ generated from the Gaussian Orthogonal Ensemble (GOE) mehta , they have zero mean and variance $\langle H_{n,n^{\prime}}^{2}\rangle=(1+\delta_{n,n^{\prime}})/(4(N+1))$. The averaged density of states is described by the the semi-circle law $dm/dE=\frac{2N}{\pi}\sqrt{1-E^{2}}$ with typical eigenvalues in the interval $E_{m}\in[-1,1]$ (we use dimensionless units with $\hbar=1$). Here, $\beta$ is a dimensionless constant characterizing the nonlinear interaction strength in the original basis $n$. In a first approximation we can consider that the eigenenergies $E_{m}$ are changing linearly with $m$ in the range [1,1][-1,1], however, keeping in mind that at the spectrum boundaries the density of states drops significantly.
@@
corresponding to the equipartition of energies $\langle(E_{m}-\mu)|C_{m}|^{2}\rangle=(E_{m}-\mu)\rho_{m}=T$ where $T$ is the system temperature, $\mu(T)$ is the chemical potential dependent on temperature. These two parameters are determined from the total norm [^1] For simplicity of notation we use the notation norm for the quantity $\kappa$ even though it is a certain time averaged norm $|C_{m}(t)|^{2}$ of the quantum state with amplitudes $C_{m}(t)$. Actually, $\kappa$ is mathematically the 1-norm of the vector with coefficients $\rho_{m}$. Furthermore, we denote as norm distribution the dependence of $\rho_{m}$ on energies $E_{m}$ in particular for the cases where $\kappa\neq 1$. $\kappa\equiv\sum_{m}\rho_{m}=1$ and the energy $E$, which are conserved integrals of motion, by the implicit equations $\kappa=\sum_{m}\rho_{EQ}(E_{m})=1$ and $\sum_{m}E_{m}\rho_{EQ}(E_{m})=E$ (for $E$ we assume the case of a weak or moderate nonlinearity which provides only a weak contribution to the total energy). The entropy $S$ of the system is given by the usual relation landau : $S=-\sum_{m}\rho_{m}\ln\rho_{m}$ with the implicit theoretical dependencies on temperature $E(T)$, $S(T)$, $\mu(T)$ (see details in rmtprl ).
@@
$i{{\partial C_{m}}\over{\partial{t}}}$ $=$ $E_{m}C_{m}+(\gamma_{m}-\sigma_{m}|C_{m}|^{2})$
@@
In (II), non-zero values of $\gamma_{m}>0$, $\sigma_{m}>0$ correspond to pumping modes or $\gamma_{m}<0$, $\sigma_{m}=0$ to dissipation modes. To obtain the energy flow of a direct cascade from low to high energy modes $m$ we choose for pumping $\gamma_{m}=\gamma>0$ for the 4 lowest energy modes at $m=1,2,3,4$ with corresponding saturation coefficients $\sigma_{m}=\sigma>0$ and for dissipation $\gamma_{m}=-\gamma<0,\sigma_{m}=0$ for the 4 highest energy modes with m=N,N1,N=2,N3m=N,N-1,N=2,N-3. For all other $m$ values we chose $\gamma_{m}=\sigma_{m}=0$. Here $\gamma$ and $\sigma$ are two parameters of our model and in most cases we choose $\gamma=\sigma=0.01$.
@@
$C_{m}(t)=\frac{C_{m}(0)\,e^{-iE_{m}t}}{\sqrt{(D_{m}+(1-D_{m})e^{-2\gamma_{m}t}}}\ ,\ D_{m}\equiv|C_{m}(0)|^{2}\frac{\sigma_{m}}{\gamma_{m}},$ (4)
@@ IV Numerical results for RMM of KZT
Figure 1: Distribution $\rho_{m}$ versus $E_{m}+1$ in a double logarithmic representation for $N=128$ (red ++) and $N=256$ (blue *), $\gamma=\sigma=0.01$, $\beta=1$, pumping at $m=1,\ldots,4$ and absorption at $m=N-3,\ldots,N$. The values of $\rho_{m}$ have been obtained from the time average $\rho_{m}(t)=\langle|C_{m}(\tau)|^{2}\rangle$ for $t/2\leq\tau\leq t$ with $t=2^{22}$. The initial condition corresponds to uniform random values $C_{m}(0)$ for $m=1,\ldots,8$ and $C_{m}(0)=0$ for $m>8$ with initial squared norm $\sum_{m}|C_{m}(0)|^{2}=0.01$. Data points with $E_{m}+1<0.09$ have been artificially moved to $E_{m}+1=0.09$. The black straight lines show the power law fit $\rho_{m}=A\,(1+E_{m})^{-s_{0}}$ using the fit range $0.2\leq 1+E_{m}\leq 1.8$ with $A=(9.52\pm 0.10)\times 10^{-3}$, $s_{0}=1.48\pm 0.02$ (for $N=128$) and $A=(2.69\pm 0.01)\times 10^{-3}$, $s_{0}=1.24\pm 0.01$ (for $N=256$). The average total norm is $\kappa=\sum_{m}\rho_{m}=6.25$ (for $N=128$) and $\kappa=5.31$ (for $N=256$). The green curves show the theoretical equipartition distribution $\rho_{EQ}(E_{m})$ of Eq. (2) obtained from a reduced spectrum 4<m<N34<m<N-3, excluding both pumping and dissipation modes with effective temperature $T=8.40\times 10^{-3}$ and chemical potential $\mu=-0.881$ (for $N=128$) and $T=2.45\times 10^{-3}$, $\mu=-0.918$ (for $N=256$; see text for details). The values of the reduced norm $\kappa_{r}$ and reduced energy $E_{r}$ used in Eq. (8) are $\kappa_{r}=2.489$, $E_{r}=-1.185$ (for $N=128$) and $\kappa_{r}=1.417$, $E_{r}=-0.6930$ (for $N=256$).
@@ IV.1 Direct cascade of KZT
One can also ask the question in how far the numerical results for Fig. 1 correspond to the statistical equipartition distribution (2). Obviously, the numerical values of $\rho_{m}$ in Fig. 1 are significantly enhanced (reduced) for pumping (dissipation) modes $m$ if compared to $\rho_{m}$ for other close $m$ non-pumping (non-dissipation) modes and also from a theoretical view point it is not reasonable to expect that Eq. (2) is valid for the full spectrum including pumping and dissipation modes. However, one may surmise that the long time ($t=2^{22}$) steady state distribution for the reduced spectrum with 4<m<N44<m<N-4, excluding pumping and dissipation modes, satisfies approximately the statistical equipartition distribution (2). To verify this point we have determined $T$ and $\mu$ by the two implicit equations :
@@
where the sums $\sum^{\prime}$ are taken over the reduced spectrum 4<m<N44<m<N-4 and the reduced norm $\kappa_{r}$ and energy $E_{r}$ are obtained from the numerical values of $\rho_{m}$ at $t=2^{22}$ by :
@@
The green curves in Fig. 1 show $\rho_{EQ}(E_{m})$ for the reduced spectrum of $N=128,256$, $\beta=1$, $\gamma=\sigma=0.01$ (the obtained values of $T,\mu,\kappa_{r},E_{r}$ are given in the figure caption of Fig. 1). These curves are rather close to the numerical data but for modes close to the energy boundaries there are some small but significant deviations and especially for energies in the interval $1<E_{m}<1.8$ the power law based on KZ theory (straight black lines) provide a slightly better fit than the classical equipartition distribution.
@@
Figure 3: As Fig. 1 but $N=512$ and modified values $\gamma=0.01$ (red ++), $\gamma=0.005$ (green $\times$ and shifted down by a factor of 10 for better visibility) and $\gamma=0.0001$ (blue *). (Values for $\beta=1$, $\sigma=0.01$, $t=2^{22}$, modes for pumping/absorption and initial condition are unchanged with respect to Fig. 1.) The straight lines show the power law fit $\rho_{m}=A\,(1+E_{m})^{-s_{0}}$ using the fit range $0.2\leq 1+E_{m}\leq 1.8$ with $A=(1.584\pm 0.003)\times 10^{-3}$, $s_{0}=1.130\pm 0.004$ (for $\gamma=0.01$), $A=(1.343\pm 0.004)\times 10^{-3}$, $s_{0}=1.117\pm 0.005$ (for $\gamma=0.005$) and $A=(1.82\pm 0.03)\times 10^{-12}$, $s_{0}=2.17\pm 0.03$ (for $\gamma=0.0001$) The average total norm is $\kappa=5.99$ (for $\gamma=0.01$), $3.58$ (for $\gamma=0.005$) and $0.0455$ (for $\gamma=0.0001$).
We show in Fig. 3 the steady-state norm distribution $\rho_{m}$ for three values of pumping $\gamma=0.01;0.005;0.0001$ and the same value $\sigma=0.01$. For two largest $\gamma$ values the pumping is sufficiently strong (strong wind on a sea) and the steady-state distribution is rather close to the KZ theory. It is important to note that this steady-state is practically independent of the initial values of amplitudes $C_{m}(t=0)$ (for a given fixed realization of the RM $\hat{H}$). We have verified this point by choosing different random realizations of the initial condition (for the random initial values $C_{m}(0)$ for $m\leq 8$ with $\kappa(t=0)=0.01$) and also localized initial conditions $C_{m_{0}}(0)=1$ for a specific mode value $m_{0}$ (e.g. with $m_{0}=1$ or $m_{0}=N/2$ etc.) and $C_{m}(0)=0$ for $m\neq m_{0}$ (typical initial condition chosen in rmtprl with $\kappa(t=0)=1$). In all cases and for $N\leq 512$, we obtain essentially the same steady-state distribution at $t=2^{22}$ with average relative differences $\langle\delta\rho_{m}/\rho_{m}\rangle$ of the order of $(1-2)\times 10^{-2}$ (for $N=128,256$) or $(5-10)\times 10^{-2}$ (for $N=512$). This shows that there a sigle global chaotic attractor with describes a turbulent flow in the system.
@@
Figure 5: As Fig. 1 for $N=1024$ and $\gamma=0.01$ (red ++) or $\gamma=0.02$ (blue *) (same other parameters). The straight lines show the power law fit $\rho_{m}=A\,(1+E_{m})^{-s_{0}}$ using the fit range $0.2\leq 1+E_{m}\leq 1.8$ with $A=(7.64\pm 0.04)\times 10^{-6}$, $s_{0}=2.43\pm 0.01$ (for $\gamma=0.01$) and $A=(2.713\pm 0.005)\times 10^{-4}$, $s_{0}=1.134\pm 0.003$. The total norm is $\kappa=4.15$ (for $\gamma=0.01$) and $9.07$ (for $\gamma=0.02$).
@@ IV.2 Regime of inverse cascade
According to KZ theory zakharovbook there should also be an inverse turbulent cascade of norm flow with the algebraic exponent $x_{0}=2/3$ (7). There have been experiments (see e.g. luka ) and numerical simulations (see e.g. shrira ) of realizations of such an inverse cascade. We try to realize the regime of an inverse cascade (for $\beta=1$, $\gamma=\sigma=0.01$, $N=512$) by performing $\gamma$-pumping in the same way as before but placing it either in the middle of the energy band $E_{m}$ (4 modes near $m_{0}=N/2$) or in the vicinity of maximal $E_{m}$ energies (4 modes near $m_{0}=N-40$). The absorption is done at both energy boundaries at 8 eigenmodes with $m=1,\ldots,4$ and $m=N-3,\ldots,N$ (using $\gamma_{m}=-\gamma<0$ and $\sigma_{m}=0$). The initial state is always composed of 8 eigenmodes placed around $m_{0}$ with random amplitudes and the total initial norm $\kappa(t=0)\approx 0.01$.
@@
Figure 7: Time dependence of the (non-averaged) norm $\sum_{m}|C_{m}(t)|^{2}$ at selected time values for $N=512$, $\beta=1$, $\gamma=\sigma=0.01$. The pumping/absorption parameters are as in Fig. 2 for data points (a) (red ++) or in (b) panel of Fig. 6 for data points (b) (blue *). Note that the total norm for times $\tau$ in the interval $2^{21}\leq\tau\leq 2^{22}$ is $\kappa=5.99$ (a) and $\kappa=48.52$ (b) (see also captions of Figs. 2 and 5).

### v2

@@ I Introduction
The Kolmogorov concept of turbulence kolm41 ; obukhov assumes emergence of energy flow through an inertial interval from large spacial scales, with an external pumping, to small scales where energy is absorbed by dissipation. The scaling arguments lead to appearance of a power law energy distribution over wave modes for hydrodynamics turbulence kolm41 ; obukhov . This concept was generalized and extended to weak wave turbulence, based on diagrammatic techniques and the kinetic equation, indeed showing the emergence of power law distributions for various types of weakly interacting nonlinear waves filonenko ; zakharovbook ; nazarenkobook ; luka ; galtier . This theory became known as Kolmogorov-Zakharov (KZ) turbulence (or spectra) nazarenkobook ; luka ; galtier . In spite of various successful confirmations of this theory in experiments and numerical simulations(see e.g. nazarenkobook ; luka ; galtier ) it is still based on a fundamental hypothesis directly stated in the seminal work of Zhakharov and Filonenko filonenko : “In the theory of weak turbulence nonlinearity of waves is assumed to be small; this enables us, using the hypothesis of the random nature of the phase of individual waves, to obtain the kinetic equation for the mean square of the wave amplitudes”. Nevertheless, the dynamical equations for waves do not involve Random Phase Approximation (RPA) and hence, the validity for the whole concept of energy flow from large to small scales remains open.
Indeed, a flow through an inertial interval is Hamiltonian and it is well known that in nonlinear systems with weak nonlinearity the Kolmogorov-Arnold-Moser (KAM) theory guarantees that the main part of the system phase space remains integrable and non-chaotic in the limit of very weak nonlinearity (see e.g. arnold ; sinai ; note that KAM theory is valid in absence of exact resonances as discussed below). More physical analysis of nonlinear dynamical Hamiltonian systems also shows the existence of a chaos border below which the phase space contains mainly an integrable dynamics opposite to the turbulent one chirikov ; lichtenberg .
@@ II Model description
[…]
@@
corresponding to the equipartition of energies $\langle(E_{m}-\mu)|C_{m}|^{2}\rangle=(E_{m}-\mu)\rho_{m}=T$ where $T$ is the system temperature, $\mu(T)$ is the chemical potential dependent on temperature. These two parameters are determined from the total norm111 For simplicity of notation we use the notation norm for the quantity $\kappa$ even though it is a certain time averaged norm $|C_{m}(t)|^{2}$ of the quantum state with amplitudes $C_{m}(t)$. Actually, $\kappa$ is mathematically the 1-norm of the vector with coefficients $\rho_{m}$. Furthermore, we denote as norm distribution the dependence of $\rho_{m}$ on energies $E_{m}$ in particular for the cases where $\kappa\neq 1$. $\kappa\equiv\sum_{m}\rho_{m}=1$ and the energy $E$, which are conserved integrals of motion, by the implicit equations $\kappa=\sum_{m}\rho_{EQ}(E_{m})=1$ and $\sum_{m}E_{m}\rho_{EQ}(E_{m})=E$ (for $E$ we assume the case of a weak or moderate nonlinearity which provides only a weak contribution to the total energy). The entropy $S$ of the system is given by the usual relation landau : $S=-\sum_{m}\rho_{m}\ln\rho_{m}$ with the implicit theoretical dependencies on temperature $E(T)$, $S(T)$, $\mu(T)$ (see details in rmtprl ). We note that a random matrix model similar to those considered in rmtprl and in (1) was considered in prxshapiro but a detailed study of dynamical thermalization was not presented there.
@@
$i{{\partial C_{m}}\over{\partial{t}}}$ $=$ $E_{m}C_{m}+i(\gamma_{m}-\sigma_{m}|C_{m}|^{2})C_{m}$
@@
In (II), non-zero values of $\gamma_{m}>0$, $\sigma_{m}>0$ correspond to pumping modes or $\gamma_{m}<0$, $\sigma_{m}=0$ to dissipation modes. To obtain the energy flow of a direct cascade from low to high energy modes $m$ we choose for pumping $\gamma_{m}=\gamma>0$ for the 4 lowest energy modes at $m=1,2,3,4$ with corresponding saturation coefficients $\sigma_{m}=\sigma>0$ and for dissipation $\gamma_{m}=-\gamma<0,\sigma_{m}=0$ for the 4 highest energy modes with $m=N,N-1,N=2,N-3$. For all other $m$ values we chose $\gamma_{m}=\sigma_{m}=0$. Here $\gamma$ and $\sigma$ are two parameters of our model and in most cases we choose $\gamma=\sigma=0.01$.
@@
$C_{m}(t)=\frac{C_{m}(0)\,e^{-iE_{m}t}}{\sqrt{D_{m}+(1-D_{m})e^{-2\gamma_{m}t}}}\ ,\ D_{m}\equiv|C_{m}(0)|^{2}\frac{\sigma_{m}}{\gamma_{m}},$ (4)
@@ IV Numerical results for RMM of KZT
Figure 1: Distribution $\rho_{m}$ versus $E_{m}+1$ in a double logarithmic representation for $N=128$ (red $+$) and $N=256$ (blue *), $\gamma=\sigma=0.01$, $\beta=1$, pumping at $m=1,\ldots,4$ and absorption at $m=N-3,\ldots,N$. The values of $\rho_{m}$ have been obtained from the time average $\rho_{m}(t)=\langle|C_{m}(\tau)|^{2}\rangle$ for $t/2\leq\tau\leq t$ with $t=2^{22}$. The initial condition corresponds to uniform random values $C_{m}(0)$ for $m=1,\ldots,8$ and $C_{m}(0)=0$ for $m>8$ with initial squared norm $\sum_{m}|C_{m}(0)|^{2}=0.01$. Data points with $E_{m}+1<0.09$ have been artificially moved to $E_{m}+1=0.09$. The black straight lines show the power law fit $\rho_{m}=A\,(1+E_{m})^{-s_{0}}$ using the fit range $0.2\leq 1+E_{m}\leq 1.8$ with $A=(9.52\pm 0.10)\times 10^{-3}$, $s_{0}=1.48\pm 0.02$ (for $N=128$) and $A=(2.69\pm 0.01)\times 10^{-3}$, $s_{0}=1.24\pm 0.01$ (for $N=256$). The average total norm is $\kappa=\sum_{m}\rho_{m}=6.25$ (for $N=128$) and $\kappa=5.31$ (for $N=256$). The green curves show the theoretical equipartition distribution $\rho_{EQ}(E_{m})$ of Eq. (2) obtained from a reduced spectrum $34<m<N-$ 3, excluding both pumping and dissipation modes with effective temperature $T=8.40\times 10^{-3}$ and chemical potential $\mu=-0.881$ (for $N=128$) and $T=2.45\times 10^{-3}$, $\mu=-0.918$ (for $N=256$; see text for details). The values of the reduced norm $\kappa_{r}$ and reduced energy $E_{r}$ used in Eq. (8) are $\kappa_{r}=2.489$, $E_{r}=-1.185$ (for $N=128$) and $\kappa_{r}=1.417$, $E_{r}=-0.6930$ (for $N=256$).
@@ IV.1 Direct cascade of KZT
One can also ask the question in how far the numerical results for Fig. 1 correspond to the statistical equipartition distribution (2). Obviously, the numerical values of $\rho_{m}$ in Fig. 1 are significantly enhanced (reduced) for pumping (dissipation) modes $m$ if compared to $\rho_{m}$ for other close $m$ non-pumping (non-dissipation) modes and also from a theoretical view point it is not reasonable to expect that Eq. (2) is valid for the full spectrum including pumping and dissipation modes. However, one may surmise that the long time ($t=2^{22}$) steady state distribution for the reduced spectrum with $44<m<N-$ 4, excluding pumping and dissipation modes, satisfies approximately the statistical equipartition distribution (2). To verify this point we have determined $T$ and $\mu$ by the two implicit equations :
@@
where the sums $\sum^{\prime}$ are taken over the reduced spectrum $44<m<N-$ 4 and the reduced norm $\kappa_{r}$ and energy $E_{r}$ are obtained from the numerical values of $\rho_{m}$ at $t=2^{22}$ by :
@@
The green curves in Fig. 1 show $\rho_{EQ}(E_{m})$ for the reduced spectrum of $N=128,256$, $\beta=1$, $\gamma=\sigma=0.01$ (the obtained values of $T,\mu,\kappa_{r},E_{r}$ are given in the figure caption of Fig. 1). These curves are rather close to the numerical data but for modes close to the energy boundaries there are some small but significant deviations and especially for energies in the interval $1.81<E_{m}<$ 1.8 the power law based on KZ theory (straight black lines) provide a slightly better fit than the classical equipartition distribution.
@@
Figure 3: As Fig. 1 but $N=512$ and modified values $\gamma=0.01$ (red $+$), $\gamma=0.005$ (green $\times$ and shifted down by a factor of 10 for better visibility) and $\gamma=0.0001$ (blue *). (Values for $\beta=1$, $\sigma=0.01$, $t=2^{22}$, modes for pumping/absorption and initial condition are unchanged with respect to Fig. 1.) The straight lines show the power law fit $\rho_{m}=A\,(1+E_{m})^{-s_{0}}$ using the fit range $0.2\leq 1+E_{m}\leq 1.8$ with $A=(1.584\pm 0.003)\times 10^{-3}$, $s_{0}=1.130\pm 0.004$ (for $\gamma=0.01$), $A=(1.343\pm 0.004)\times 10^{-3}$, $s_{0}=1.117\pm 0.005$ (for $\gamma=0.005$) and $A=(1.82\pm 0.03)\times 10^{-12}$, $s_{0}=2.17\pm 0.03$ (for $\gamma=0.0001$) The average total norm is $\kappa=5.99$ (for $\gamma=0.01$), $3.58$ (for $\gamma=0.005$) and $0.0455$ (for $\gamma=0.0001$).
We show in Fig. 3 the steady-state norm distribution $\rho_{m}$ for three values of pumping $\gamma=0.01;0.005;0.0001$ and the same value $\sigma=0.01$. For two largest $\gamma$ values the pumping is sufficiently strong and the steady-state distribution is rather close to the KZ theory. It is important to note that this steady-state is practically independent of the initial values of amplitudes $C_{m}(t=0)$ (for a given fixed realization of the RM $\hat{H}$). We have verified this point by choosing different random realizations of the initial condition (for the random initial values $C_{m}(0)$ for $m\leq 8$ with $\kappa(t=0)=0.01$) and also localized initial conditions $C_{m_{0}}(0)=1$ for a specific mode value $m_{0}$ (e.g. with $m_{0}=1$ or $m_{0}=N/2$ etc.) and $C_{m}(0)=0$ for $m\neq m_{0}$ (typical initial condition chosen in rmtprl with $\kappa(t=0)=1$). In all cases and for $N\leq 512$, we obtain essentially the same steady-state distribution at $t=2^{22}$ with average relative differences $\langle\delta\rho_{m}/\rho_{m}\rangle$ of the order of $(1-2)\times 10^{-2}$ (for $N=128,256$) or $(5-10)\times 10^{-2}$ (for $N=512$). This shows that there a sigle global chaotic attractor with describes a turbulent flow in the system.
@@
Figure 5: As Fig. 1 for $N=1024$ and $\gamma=0.01$ (red $+$) or $\gamma=0.02$ (blue *) (same other parameters). The straight lines show the power law fit $\rho_{m}=A\,(1+E_{m})^{-s_{0}}$ using the fit range $0.2\leq 1+E_{m}\leq 1.8$ with $A=(7.64\pm 0.04)\times 10^{-6}$, $s_{0}=2.43\pm 0.01$ (for $\gamma=0.01$) and $A=(2.713\pm 0.005)\times 10^{-4}$, $s_{0}=1.134\pm 0.003$. The total norm is $\kappa=4.15$ (for $\gamma=0.01$) and $9.07$ (for $\gamma=0.02$).
@@ IV.2 Regime of inverse cascade
According to KZ theory zakharovbook there should also be an inverse turbulent cascade of norm flow with the algebraic exponent $x_{0}=2/3$ (7). There have been experiments (see e.g. luka ) and numerical simulations (see e.g. shrira ; nazar2023 ) of realizations of such an inverse cascade. We try to realize the regime of an inverse cascade (for $\beta=1$, $\gamma=\sigma=0.01$, $N=512$) by performing $\gamma$-pumping in the same way as before but placing it either in the middle of the energy band $E_{m}$ (4 modes near $m_{0}=N/2$) or in the vicinity of maximal $E_{m}$ energies (4 modes near $m_{0}=N-40$). The absorption is done at both energy boundaries at 8 eigenmodes with $m=1,\ldots,4$ and $m=N-3,\ldots,N$ (using $\gamma_{m}=-\gamma<0$ and $\sigma_{m}=0$). The initial state is always composed of 8 eigenmodes placed around $m_{0}$ with random amplitudes and the total initial norm $\kappa(t=0)\approx 0.01$.
@@
Figure 7: Time dependence of the (non-averaged) norm $\sum_{m}|C_{m}(t)|^{2}$ at selected time values for $N=512$, $\beta=1$, $\gamma=\sigma=0.01$. The pumping/absorption parameters are as in Fig. 2 for data points (a) (red $+$) or in (b) panel of Fig. 6 for data points (b) (blue *). Note that the total norm for times $\tau$ in the interval $2^{21}\leq\tau\leq 2^{22}$ is $\kappa=5.99$ (a) and $\kappa=48.52$ (b) (see also captions of Figs. 2 and 5).
@@
Here we consider finite size systems with random matrix couplings between modes. Such a situation corresponds to multimode fibers with a cross-section generating chaotic ray dynamics corresponding to a regime of quantum chaos (e.g. D-shape cross-section as discussed in rmtprl ); such fibers may have up to thousand of modes). Such systems are rather different from quasi-infinite sizes used to study weak turbulence of waves in large systems nazarenkobook ; shrira ; nazar2023 (e.g. about billion modes in nazar2023 ). It was shown that at large sizes and a specific pumping the theoretical inverse cascade can be realized nazar2023 . We do not exclude that for some specific pumping and very large random matrix sizes one can find regimes with theoretical inverse cascade. But for moderate sizes of about thousand, typical for fibers, we tried various typical forms of pumping that were always leading to a regime of strong nonlinearity. Thus we conclude that the theoretical regime of inverse cascade is hardly reachable for finite size systems with random matrix interactions between modes, e.g. chaotic fibers. Also we think that for fibers it is rather difficult to pump only very high modes.